An infinite plane of charge has surface charge density 5.8 μC/m2. How far apart are the equipotential surfaces whose potentials differ by 100 eBook
Answer #1

Solution)

Given,

Surface charge density,σ = 5.8 *10^-6 C/m^2

We know, according to Gauss's law,

For an infinite plane sheet of charge:

E = σ / (2 * ε_0)

here ε_0 = 8.85 * 10 ^ -12 C^2/N.m^2 (permeability)

E = 327683.6 N/C = constant

we know that

ΔV = E * Δd = 100V

Δd = ΔV/E

Δd =0.000305 m (Ans)

==========

Good luck!:)